OpenStudy (ashy98):

Part A: Solve A = (x + 28) for x. Part B: Determine the value of x when A = 135. Part C: Solve -np - 40 > 10 for n. Show your work.

OpenStudy (anonymous):

for Part there isn'r much to solve, just make x the subject so for A=x+28 x=A-28

OpenStudy (anonymous):

for part B if you insert the value of A into A=x+28 so it should be 135=x+28 make x the subject x=135-25 so now solve for x